∞Σn=1 sin(1/n) is a convergent series.


True or false with full explanation

1
Expert's answer
2021-11-15T15:55:20-0500

It is known that the harmonic series "\sum\limits_{n = 1}^\infty {\frac{1}{n}}" diverges

Since

"\mathop {\lim }\limits_{n \to \infty } \frac{{\sin \frac{1}{n}}}{{\frac{1}{n}}} = \mathop {\lim }\limits_{\frac{1}{n} \to 0} \frac{{\sin \frac{1}{n}}}{{\frac{1}{n}}} = 1" and "0 < 1 < \infty" than, by Limit comparison test, the series "\sum\limits_{n = 1}^\infty {\sin \frac{1}{n}}" also diverges.

Answer: False